### v1

@@ Abstract
To address the increasing complexity and frequency of cybersecurity incidents emphasized by the recent cybersecurity threat reports with over 10 billion instances, cyber threat intelligence (CTI) plays a critical role in the modern cybersecurity landscape by offering the insights required to understand and combat the constantly evolving nature of cyber threats. Inspired by the powerful capability of large language models (LLMs) in handling complex tasks, in this paper, we introduce a framework to benchmark, elicit, and improve cybersecurity incident analysis and response abilities in LLMs for Security Events (called SEvenLLM). Specifically, we create a high-quality bilingual instruction corpus by crawling cybersecurity raw text from cybersecurity websites to overcome the lack of effective data for information extraction. Then, we design a pipeline to auto-select tasks from the tasks pool and convert the raw text into supervised corpora comprised of question and response. The instruction dataset SEvenLLM-Instruct is used to train cybersecurity LLMs with the multi-task learning objective (27 well-designed tasks) for augmenting the analysis of cybersecurity events. Extensive experiments in our curated benchmark (SEvenLLM-Bench) demonstrate that SEvenLLM performs more sophisticated threat analysis and fortifies defenses against the evolving landscape of cyber threats [^1] https://github.com/CSJianYang/SEevenLLM.
@@ 1 Introduction
In this paper, we introduce a framework to benchmark, elicit, and improve cybersecurity incident analysis and response abilities in LLMs for Security Events (called SEvenLLM). In Figure 1, the fine-tuned SEvenLLM with fewer parameters can get a more complete and professional response compared to the strong common LLM GPT-3.5, emphasizing the importance of domain-specific LLM for CTI. Specifically, we address the scarcity of high-quality, task-specific datasets for security event analysis by curating an extensive bilingual (English and Chinese) corpus of cybersecurity incident reports. To construct the supervised question-answer pair, we first use LLM to generate the candidate tasks, and then human experts correct the tasks to form a task pool. Given the cybersecurity raw text, we prompt the LLM to select the proper task from the task pool and then simultaneously generate the query and corresponding response of the chosen task, where we use Select-Instruct to generate the instruction corpus SEvenLLM-Instruct. Open-source LLMs tailored for cyber threat intelligence based on different base models are fine-tuned on the SEvenLLM-Instruct with multi-task learning. To effectively evaluate SEvenLLM for cyber threat intelligence, we construct an evaluation benchmark SEvenLLM-Bench to comprehensively evaluate the performance of LLMs for CTI, thereby bridging the gap between the common field and the cybersecurity field in understanding and generation.
@@
  • To fill in the gaps in the evaluation for cyber threat intelligence, we construct an evaluation benchmark SEvenLLM-Bench comprised of multiple-choice questions and generation questions for assessing the performance of LLMs within the context of cybersecurity.
Figure 2: Overview of SEvenLLM. By crawling different formats of files from the Internet, we collect bilingual (English and Chinese) collection of cybersecurity incident reports. First, we adopt LLMs to produce potential tasks and refine them to create a task pool. Given raw cybersecurity texts, we prompt the LLM to select a proper task from the task pool and then generate the query and its answer (Select-Instruct). The open-source LLMs are further fine-tuned on SEvenLLM-Instruct with multi-task learning objectives tailored for CTI. A curated CTI evaluation benchmark SEvenLLM-Bench is created to compare SEvenLLM with other baselines.
@@ 2.2 Benchmark Construction
The analysis of cybersecurity incidents is categorized into understanding and generation tasks, encompassing 28 subcategories in Figure 2. The understanding task in cybersecurity incident analysis involves transforming unstructured data into structured data by extracting entities and relationships from cybersecurity incidents, aiming at acquiring well-defined knowledge such as identifying advanced persistent threat (APT) groups, attack techniques, and the relationships between APT groups and techniques. For cybersecurity professionals, the generation tasks (e.g. generating summaries and suggesting mitigation measures) provide straightforward knowledge for the general analyst and are used for data fusion and information compression of cybersecurity incident content. In Figure 2, we randomly select the raw text feed into the LLM to generate the task name and then correct them with the human experts to design 13 understanding tasks and 15 generation tasks. Given the cybersecurity-related raw text, we prompt the LLM to generate the supervised instruction corpora SEvenLLM-Instruct.
@@
We use multiple-choice questions and generation questions are used for to evaluate the performance of SEvenLLM in CTI. In Table 1, SEvenLLM-Instruct contains nearly 85K samples and SEvenLLM-Bench has 1200 test samples. Especially, for the multiple-choice question (MCQ), $364/1$ denotes that the length of the prompt is $364$ and the length of the output is $1$ (A, B, C, or D).
@@
Table 1: Statistics of SEvenLLM-Bench. The training data contains two types of questions: multiple-choice questions (MCQ) and generation questions (QA). The supervised data derives from high-quality raw texts (Seed) collected from the Internet. Here, we calculate the length of Chinese at the token level and the length of English by splitting space tokens.
2.3 SEvenLLM
Given a cybersecurity raw text $c\in L_{all}=\{D^{r}_{L_{i}}\}_{i=1}^{m}$ ($L_{i}$ is the language of $c$ and $L_{all}$ contains Chinese and English in our work), we prompt the LLM $\mathcal{M}$ to select the proper task $T_{k}$, where $T_{k}\in T_{all}=\{T_{j}\}_{j=1}^{n}$ and the task pool $T_{all}$ contains $13$ understanding tasks and $15$ generation tasks, as shown in Table 2. Conditioned on {Pre-defined tasks} and {Cybersecurity-related Raw Text}, we use the prompt “Based on the content and tasks of network security events, please generate suitable task category and the reason” in Figure 2 to generate the target task $T_{k}$ with the chosen reason. Then, we use the raw text $c$ and the chosen task $T_{k}$ to generate the query, the corresponding response, and the thought $q_{k},a_{k},t_{k}$ to construct the supervised fine-tuned corpora $D_{all}=\{D_{i}\}_{i=1}^{m}$, where $D_{all}$ is the supervised instruction corpus containing $m$ tasks and $(q_{k},a_{k})\in D^{s}_{all}$. In Figure 2, we feed the prompt “{Cybersecurity-related Raw Text}. {Task Definition}. Based on the content and the task of network security events, please generate structured data including input, instruction, thought, and output” into the LLM $\mathcal{M}$ to generate the supervised pair $q_{k},a_{k},t_{k}$. The CoT generation process can be described as:
@@ 2.4 Multitask Instruction Tuning
(1) Key Entity Recognition Identify the main entity information in the text, such as attacker organization, victim type, main person, vulnerability number, email address, virtual account, IP address, and other information.
(2) Main Relation Extraction Relationships between entities such as attacker, victim, attack method, and attack target. Through relationship extraction, connections between entities can be established to help cybersecurity experts better understand the content and context of threat intelligence.
@@
(4) Malware Feature Extraction Extract the malware features mentioned in the text, such as file characteristics, behavioral patterns, means of exploitation, and harm caused.
@@
(11) Vulnerability Information Extraction Characterize the attacker’s place of origin, ip, language, unique fingerprints during the attack, and attack behavior and confrontation ideas/
(12) Attacker Information Extraction Identify the main information of the vulnerability described in the text, the vulnerability number, and the way the vulnerability is exploited.
(13) Attack Target Intelligence Gathering Characterize the attacker’s place of origin, ip, language, unique fingerprints during the attack, and attack behavior and confrontation ideas.
@@
(2) Attack Means Analysis Analysis of the means and specific methods used in attacks during cybersecurity incidents.
@@
(4) Correlation Analysis Analyze the scope and impact of security incidents.
@@
(10) Behavioral Pattern Analysis Research the behavioral pattern characteristics of attackers in network security incidents, including attack timing, target selection, etc.
@@
Qwen1.5-Chat Bai et al. (2023) Qwen 7B 78.2 79.5 79.5 77.5 78.9 78.5 78.7
Qwen1.5-Chat Bai et al. (2023) Qwen 14B 82.1 79.6 79.7 77.7 80.9 78.6 79.8
SEvenLLM Llama 7B 82.1 79.7 80.9 77.9 81.5 78.8 80.2
SEvenLLM + CoT Llama 7B 82.3 79.7 79.6 77.5 81.0 78.6 79.8
SEvenLLM Llama 13B 83.7 80.3 80.9 78.0 82.3 79.1 80.7
SEvenLLM + CoT Llama 13B 82.7 80.1 80.3 78.3 81.5 79.2 80.4
SEvenLLM Qwen 7B 84.1 80.7 81.3 78.7 82.7 79.7 81.2
SEvenLLM Qwen 14B 82.6 79.7 80.6 77.9 81.6 78.8 80.2
@@ 3.1 Evaluation
Using the multilingual sentence-transformers model [^5] https://huggingface.co/sentence-transformers/paraphrase-multilingual-MiniLM-L12-v2 Reimers and Gurevych (2019), we map the text to a 384-dimensional dense vector space. This can be used for tasks such as clustering or semantic search, allowing for the evaluation of whether the values in a test dictionary can score from a semantic understanding perspective. This approach is more aligned with the open-ended nature of cyber security event analysis. Due to its support for various lengths and both Chinese and English languages, this model is applied to both extractive and generative tasks.
@@ 4 Main Results
In Table 3, compared to the results of models that have not been trained with SEvenLLM-Instruct, it is evident that training with the datasets provided in this work significantly improves performance. Additionally, it is observed that while higher-capacity models show improvements in multiple-choice questions, they can occasionally perform worse in fuzzy testing and semantic similarity matching. This reinforces the previous analysis and highlights that within the sub-20B parameter range, larger models may not necessarily include more content in the domain of cybersecurity event analysis within their pre-training corpora. Considering that cybersecurity data is challenging to collect and highly specialized, this underscores the importance of enhancing training materials in this field with datasets like those provided in this paper. It has also been observed that base models can perform correct analysis but still produce incorrect answers, indicating that the full potential of an AI model requires not only general training but also alignment with specialized corpora to yield more accurate outcomes.
Rouge-L&Semantic Evaluation
In Table 4 and Table 5, numerous repetitive outputs are observed in Qwen-14B-Chat, which is a tendency in language model decoding to increase the probability of repeating previous sentences. Since the parameters are consistent with the 7B model, it is inferred that higher parameter models may require adjustments to the repetition_penalty and temperature settings. Furthermore, using Llama-2-13b with CoT, where the model is guided to reason before outputting a result, did not outperform the non-CoT-aided 13B model either in fuzzy matching or similarity matching, and only performed comparably in multiple-choice questions without constructing CoT. This supports the analysis that within the 7B-13B parameter range, the introduction of guided thinking may lead to more rigid model behavior and does not necessarily improve performance.
@@
In Table 6, during a comparison of multiple-choice questions, it is observed that the Qwen performs less effectively than Llama2, particularly in scenarios requiring a degree of reasoning and where answers may be confusing. This includes questions where certain options included content from the original text, where Qwen had trouble identifying correctly. From tests carried out using fuzzy matching and semantic matching, the model’s performance appears to reach a level close to the testing set when it scales to around the 70,000 magnitude. However, overall it seems that the larger the training data set, the better the accuracy of the model. The LLMs with 13B and 14B parameters have been confirmed to perform better in multiple-choice question scenarios. However, there have been mixed results in terms of semantic similarity and fuzzy matching. This suggests that the enrichment of overall knowledge and reasoning capabilities with larger model sizes improves performance in multiple-choice questions, but the increased volume of information can lead to more divergent responses, which may not always align well with strict fuzzy and semantic matching tests in the test set. To corroborate this, the results of Llama-2-13B were scored using GPT-4, which showed an improvement, indicating that a larger model size results in richer descriptions, supporting the analysis.
@@ 5 Analysis
Figure 3: Evaluation results of different instruction data sizes. With the data size increasing, SEvenLLM gets better performance.
@@
To discuss the effect of the instruction tuning, we plot GPT-4 scores and Rouge-L scores with different training data sizes in Figure 3. The performance is influenced by the size of the instruction data. We randomly sample $N=$$\{1K,2K,\dots,ALL\}$ sentences from the whole corpora to fine-tune the Llama model. With the training data size increasing, the SEvenLLM gets better performance. Notably, only $1K$ supervised data sentences bring large improvement to the zero-shot cross-lingual NER, which benefits from knowledge transfer of the multilingual self-training. When the size of target annotated corpora is greater than 10K, our method gets exceptional performance.
@@
To ensure the quality of the SEvenLLM-Bench, we employ three experts to evaluate the quality of the test set and calculate the error rate in Figure 7. Manual revision includes corrections for accuracy and reasonableness, inspecting the result content to ensure it meets the requirements of the instructions, is based on prior knowledge from the input, and involves removing redundant information and hallucinations, as well as adding some undetected information. This includes proofreading 100 multiple-choice questions for the correctness and uniqueness of answers. We can that the total error rate of SEvenLLM-Bench is 17% generated by GPT-4 and then the experts fix these errors to ensure the correctness of the test set.
@@
Figure 4: Comparison between SEvenLLM with the GPT-3.5.

### v2

@@ Abstract
To address the increasing complexity and frequency of cybersecurity incidents emphasized by the recent cybersecurity threat reports with over 10 billion instances, cyber threat intelligence (CTI) plays a critical role in the modern cybersecurity landscape by offering the insights required to understand and combat the constantly evolving nature of cyber threats. Inspired by the powerful capability of large language models (LLMs) in handling complex tasks, in this paper, we introduce a framework to benchmark, elicit, and improve cybersecurity incident analysis and response abilities in LLMs for Security Events (called SEvenLLM). Specifically, we create a high-quality bilingual instruction corpus by crawling cybersecurity raw text from cybersecurity websites to overcome the lack of effective data for information extraction. Then, we design a pipeline to auto-select tasks from the tasks pool and convert the raw text into supervised corpora comprised of question and response. The instruction dataset SEvenLLM-Instruct is used to train cybersecurity LLMs with the multi-task learning objective (28 well-designed tasks) for augmenting the analysis of cybersecurity events. Extensive experiments in our curated benchmark (SEvenLLM-Bench) demonstrate that SEvenLLM performs more sophisticated threat analysis and fortifies defenses against the evolving landscape of cyber threats [^1] https://github.com/CSJianYang/SEevenLLM.
@@ 1 Introduction
In this paper, we introduce a framework to benchmark, elicit, and improve cybersecurity incident analysis and response abilities in LLMs for Security Events (called SEvenLLM). In Figure 1, the fine-tuned SEvenLLM with fewer parameters can get a more complete and professional response compared to the strong common LLM GPT-3.5, emphasizing the importance of domain-specific LLM for CTI. Specifically, we address the scarcity of high-quality, task-specific datasets for security event analysis by curating an extensive bilingual (English and Chinese) corpus of cybersecurity incident reports. To construct the supervised question-answer pair, we first use the LLM (GPT-4) to generate the candidate tasks, and then human experts correct the tasks based on threat intelligence analysis criteria and focal points proposed by security organizations and institutions, such as MITRE [^2] https://attack.mitre.org/ and OASIS CTI TC [^3] https://oasis-open.github.io/cti-documentation/, to form a task pool. Given the cybersecurity raw text, we prompt the LLM to select the proper task from the task pool and then simultaneously generate the query and corresponding response of the chosen task, where we use Select-Instruct to generate the instruction corpus SEvenLLM-Instruct. Open-source LLMs tailored for cyber threat intelligence based on different base models are fine-tuned on the SEvenLLM-Instruct with multi-task learning. To effectively evaluate SEvenLLM for cyber threat intelligence, we construct an evaluation benchmark SEvenLLM-Bench to comprehensively evaluate the performance of LLMs for CTI, thereby bridging the gap between the common field and the cybersecurity field in understanding and generation.
@@
  • To fill in the gaps in the evaluation for cyber threat intelligence, we construct an evaluation benchmark SEvenLLM-Bench comprised of multiple-choice questions and query-answer questions for assessing the performance of LLMs within the context of cybersecurity.
Figure 2: Overview of SEvenLLM. By crawling different formats of files from the Internet, we collect bilingual (English and Chinese) collection of cybersecurity incident reports. First, we adopt LLMs to produce potential tasks and refine them to create a task pool. Given raw cybersecurity texts, we use Select-Instruct select a proper task and generate the query and its answer. The open-source LLMs are further fine-tuned on SEvenLLM-Instruct with multi-task learning objectives tailored for CTI. A curated CTI evaluation benchmark SEvenLLM-Bench is created to compare SEvenLLM with other baselines.
@@ 2.2 Benchmark Construction
Select-Instruct
To improve the quality and informational value of the generated instruction-answer datasets, we have improved the self-instructWang et al. (2023a) method specifically for the task of cybersecurity incident analysis. Our approach involves two steps. In the first step, we input the raw corpus and use carefully designed prompts to enable the LLM to select a specified number and format of tasks from a task pool, producing task seeds. In the second step, we input both the task seeds and the raw corpus, requiring the LLM to extract valuable portions of the corpus based on the tasks to generate instructions and corresponding answers. During the generation of instructions and responses, we also require the LLM to provide its reasoning process and justificationJin et al. (2024). Testing results demonstrate that, using our method, out of 44,240 generated data points, there were only 24 erroneous data points and 37 empty outputs. In contrast, the self-instruct method, when used to generate instructions and answers for all tasks, produced 41,296 data points, of which 78 were erroneous and 1,182 were empty outputs.
@@
The analysis of cybersecurity incidents is categorized into understanding and generation tasks, encompassing 28 subcategories in Figure 2. The understanding task in cybersecurity incident analysis involves transforming unstructured data into structured data by extracting entities and relationships from cybersecurity incidents, aiming at acquiring well-defined knowledge such as identifying advanced persistent threat (APT) groups, attack techniques, and the relationships between APT groups and techniques. For cybersecurity professionals, the generation tasks (e.g. generating summaries and suggesting mitigation measures) provide straightforward knowledge for the general analyst and are used for data fusion and information compression of cybersecurity incident content. In Figure 2, we randomly select the raw text feed into the LLM to generate the task name and then correct them with the human experts to design 13 understanding tasks and 15 generation tasks. Given the cybersecurity-related raw text, we use Select-Instruct method to generate the supervised instruction corpora SEvenLLM-Instruct.
@@
We use multiple-choice questions and query-answer questions to evaluate the performance of SEvenLLM in CTI. In Table 1, SEvenLLM-Instruct contains nearly 90K samples and SEvenLLM-Bench has 1300 test samples. Especially, for the multiple-choice question (MCQ), $364/1$ denotes that the length of the prompt is $364$ and the length of the output is $1$ (A, B, C, or D).
@@
Table 1: Statistics of SEvenLLM-Bench. The training data contains two types of questions: multiple-choice questions (MCQ) and query-answer questions(QA). The supervised data derives from high-quality raw texts (Seed) collected from the Internet. Here, we calculate the length of Chinese at the token level and the length of English by splitting space tokens.
2.3 Large Language Model for CTI
Given a cybersecurity raw text $c\in L_{all}=\{D^{r}_{L_{i}}\}_{i=1}^{m}$ ($L_{i}$ is the language of $c$ and $L_{all}$ contains Chinese and English in our work), we prompt the LLM $\mathcal{M}$ to select the proper task $T_{k}$, where $T_{k}\in T_{all}=\{T_{j}\}_{j=1}^{n}$ and the task pool $T_{all}$ contains $13$ understanding tasks and $15$ generation tasks, as shown in Table 2. Conditioned on {Pre-defined tasks} and {Cybersecurity-related Raw Text}, we use the prompt “Based on the content and tasks of cyber security events, please generate suitable task category and the reason” in Figure 2 to generate the target task $T_{k}$ with the chosen reason. Then, we use the raw text $c$ and the chosen task $T_{k}$ to generate the query, the corresponding response, and the thought $q_{k},a_{k},t_{k}$ to construct the supervised fine-tuned corpora $D_{all}=\{D_{i}\}_{i=1}^{m}$, where $D_{all}$ is the supervised instruction corpus containing $m$ tasks and $(q_{k},a_{k})\in D^{s}_{all}$. In Figure 2, we feed the prompt “{Cybersecurity-related Raw Text}. {Task Definition}. Based on the content and the task of cyber security events, please generate structured data including input, instruction, thought, and output” into the LLM $\mathcal{M}$ to generate the supervised pair $q_{k},a_{k},t_{k}$. The CoT generation process can be described as:
@@ 2.4 Multitask Instruction Tuning
(1) Key Entity Recognition Identify the main entity information in the text, such as attacker organization, victim type, main person, the common vulnerabilities and exposures (CVE), email address, virtual account, IP address, and the indicators of compromise (IOCs).
(2) Main Relation Extraction Extract the relationships between major entities such as attacker, victim, attack method and so on. Through relationship extraction, connections between entities can be established to help cybersecurity experts better understand the content and context of threat intelligence.
@@
(4) Malware Feature Extraction Extract the malware features mentioned in the text, such as file characteristics, means of exploitation, and harm caused.
@@
(11) Vulnerability Intelligence Extraction Identify the main information of the vulnerability described in the text, the vulnerability number, and the way the vulnerability is exploited.
(12) Attacker Information Extraction Characterize the attacker’s place of origin, ip, language, unique fingerprints during the attack, and attack behavior and confrontation ideas
(13) Attack Target Intelligence Gathering Analyze the main characteristics of the attacked target, such as operating system, software, and target industry, field, geographical location, etc.
@@
(2) Attack Means Analysis Analyze the means and specific methods used in attacks during cybersecurity incidents.
@@
(4) Correlation Analysis Analyze the connections and correlated evidence between different threat intelligence reports and cybersecurity incidents.
@@
(10) Behavioral Pattern Analysis Research the behavioral pattern characteristics of attackers in cyber security incidents, including attack timing, target selection, etc.
@@
Qwen1.5-Chat Bai et al. (2023) Qwen 7B 78.2 79.5 79.5 77.5 78.9 78.5 78.7
Qwen1.5-Chat Bai et al. (2023) Qwen 14B 82.1 79.6 79.7 77.7 80.9 78.6 79.8
SEvenLLM Llama 7B 82.1 79.7 80.9 77.9 81.5 78.8 80.2
SEvenLLM + CoT Llama 7B 82.3 79.7 79.6 77.5 81.0 78.6 79.8
SEvenLLM Llama 13B 83.7 80.3 80.9 78.0 82.3 79.1 80.7
SEvenLLM + CoT Llama 13B 82.7 80.1 80.3 78.3 81.5 79.2 80.4
SEvenLLM Qwen 7B 84.1 80.7 81.3 78.7 82.7 79.7 81.2
SEvenLLM Qwen 14B 82.6 79.7 80.6 77.9 81.6 78.8 80.2
@@ 3.2 Evaluation
Using the multilingual sentence-transformers model [^7] https://huggingface.co/sentence-transformers/paraphrase-multilingual-MiniLM-L12-v2 Reimers and Gurevych (2019), we map the text to a 384-dimensional dense vector space. This can be used for tasks such as clustering or semantic search, allowing for the evaluation of whether the values in a test dictionary can score from a semantic understanding perspective. This approach is more aligned with the open-ended nature of cyber security event analysis. Due to its support for various lengths and both Chinese and English languages, this model is applied to both understanding and generation tasks.
GPT-4 Score
Multiple-choice Question Score
We design 100 multiple-choice questions for two languages and two major types of tasks to comprehensively evaluate the model’s capabilities. The output results were manually verified to ensure an effective assessment of the model’s ability to handle objective questions.
@@ 4 Main Results
Rouge-L&Semantic Evaluation
In Table 3 and Table 4, SEvenLLM based on the Llama2 and Qwen1.5 gets the best performance by fine-tuning on the instruction dataset SEvenLLM-Instruct. Further, we can observe that SEvenLLM based on the Qwen tends to get better performance in Table 3. Plus, SEvenLLM + CoT gains no improvement, since the evaluated tasks do not require the complex reason. But the reasoning text can be provided to cybersecurity staff as a reference.
@@
In Table 5, compared to the results of models that have not been trained with SEvenLLM-Instruct, it is evident that training with the datasets provided in this work significantly improves performance. Additionally, it is observed that while higher-capacity models show improvements in multiple-choice questions, they can occasionally perform worse in Rouge-L and semantic similarity matching. Considering that cybersecurity data is challenging to collect and highly specialized, this underscores the importance of enhancing training materials in this field with datasets like those provided in this paper. It has also been observed that base models can perform correct analysis but still produce incorrect answers, indicating that the full potential of an AI model requires not only general training but also alignment with specialized corpora to yield more accurate outcomes.
@@
In Table 6, during a comparison of multiple-choice questions, it is observed that the Qwen performs less effectively than Llama2, particularly in scenarios requiring a degree of reasoning and where answers may be confusing. The LLMs with 13B and 14B parameters have been confirmed to perform better in multiple-choice question scenarios. This suggests that the enrichment of overall knowledge and reasoning capabilities with larger model sizes improves performance in multiple-choice questions.
@@ 5 Analysis
Figure 3: Evaluation results of different instruction data sizes.
@@
To discuss the effect of the instruction tuning, we plot Rouge-L scores and Semantic similarity scores with different training data sizes in Figure 3. The performance is influenced by the size of the instruction data. We randomly sample $N=$$\{0K,1K,2K,\dots,ALL\}$ sentences from the whole corpora to fine-tune the Qwen model. With the training data size increasing, the SEvenLLM gets better performance. Notably, only $1K$ supervised data sentences bring large improvement to the zero-shot cross-lingual NER, which benefits from knowledge transfer of the multilingual self-training. When the size of target annotated corpora is greater than 10K, our method gets exceptional performance. From the results obtained through Rouge-L score and Semantic similarity score, the model’s performance appears to reach a level close to the testing set when it scales to around the 70,000 magnitude. However, overall it appears that the larger the training dataset, the better the accuracy of the model.
@@
To ensure the quality of the SEvenLLM-Bench, we employ three experts to evaluate the quality of the test set and calculate the error rate in Table 7. Manual revision includes corrections for accuracy and reasonableness, inspecting the result content to ensure it meets the requirements of the instructions, is based on prior knowledge from the input, and involves removing redundant information and hallucinations, as well as adding some undetected information. This includes proofreading 100 multiple-choice questions for the correctness and uniqueness of answers. We can that the total error rate of SEvenLLM-Bench is 17% generated by GPT-4 and then the experts fix these errors to ensure the correctness of the test set.
@@
Figure 4: Comparison between SEvenLLM with Llama-2-Chat.